OpenStudy (anonymous):

What is 3/8*3/8*3/8?

OpenStudy (itsbribro):

\[\frac{ 3 }{ 8 }\times \frac{ 3 }{ 8}\times \frac{ 3 }{ 8 }\] is that what you are asking?

OpenStudy (anonymous):

yes

Mehek (mehek14):

multiply \(3*3*3\) and \(8*8*8\)

OpenStudy (itsbribro):

yep ^

OpenStudy (anonymous):

Thanks!

Mehek (mehek14):

your welcome :)

OpenStudy (anonymous):

So, the answer is 27/512.

Mehek (mehek14):

yes